OpenStudy (anonymous):

Answer the following questions about the function whose derivative is given below

OpenStudy (anonymous):

\[f \prime(x)=(x-5)^{2}(x+7)\]

OpenStudy (anonymous):

a. what are the critical points of f? b. On what intervals is f increasing or decreasing? c. at what points, if any, does f assume local maximum and minimum values?

OpenStudy (agent0smith):

a) To find critical points, let f'(x) = 0 and solve for x \[\large 0 =(x-5)^{2}(x+7)\]

OpenStudy (anonymous):

x=-7, 5

OpenStudy (agent0smith):

Correct. Now, to find where f is increasing, that means f' must be positive (ie f' > 0). So check the sign of f' to the left and right of those two points... ie find if f'(-7.1) and f'(-6.9) are positive or negative. And find if f'(4.9) and f'(5.1) are positive or negative. Drawing a picture can help.  f' = 0 at -7 and 5, so find if the slope is pos/neg to the left and right of those.

OpenStudy (anonymous):

where did you get the points?

OpenStudy (anonymous):

f'(x) = 0 where there are critical points. solve for x, then plug x back into the original equation to find the y values of all the points

OpenStudy (anonymous):

that's where @agent0smith got his points

OpenStudy (anonymous):

I don't understand how

OpenStudy (agent0smith):

@clavoie ...you already posted the points... clavoie Best Response 0 x=-7, 5 which means f'(-7) = 0 and f'(5) = 0 ie the slope at x=-7 and 5 is zero. Hence why I said find if f'(-7.1) and f'(-6.9) are positive or negative. And find if f'(4.9) and f'(5.1) are positive or negative.

OpenStudy (anonymous):

where are you getting the y values?

OpenStudy (agent0smith):

What y values... you solved this \[\large 0 =(x-5)^{2}(x+7)\] for x and got x=-7 and 5... that means f' must be zero at those points, since that equation is f'.

OpenStudy (anonymous):

why is it (-7, 1)? where'd the 1 come from?

OpenStudy (agent0smith):

it's not (-7, 1). I picked a value to the left of x = -7... i picked -7.1 Reread what I wrote above: Now, to find where f is increasing, that means f' must be positive (ie f' > 0). So check the sign of f' to the left and right of those two points... ie find if f'(-7.1) and f'(-6.9) are positive or negative. And find if f'(4.9) and f'(5.1) are positive or negative. Drawing a picture can help.

OpenStudy (anonymous):

okay